OpenStudy (anonymous):

What is the standard deviation of the set of data? 21, 6, 18, 21, 20, 14, 19

OpenStudy (anonymous):

I just don't know how to find standard deviation I'm sorry ._.

OpenStudy (anonymous):

First you need to find the mean? Do you know how?

OpenStudy (anonymous):

Yup! The mean is 119.

OpenStudy (anonymous):

Oh wait I forgot a step hold on

OpenStudy (anonymous):

17

OpenStudy (anonymous):

\[mean = \frac{ 21 + 6 + 18 + 21 + 20 + 14 + 19 }{ 7 }\]

OpenStudy (anonymous):

And that's 17, right?

OpenStudy (anonymous):

Indeed :) The mean is 17

OpenStudy (anonymous):

So now, Find the Variance... Do you know how to find the variance?

OpenStudy (anonymous):

\[\sigma^2 = \frac{ (21 - 17)^2 + (6 - 17)^2 + (18 - 17)^2 + (21 - 17)^2 + (20 - 17)^2 + (14 - 17)^2 + (19 - 17)^2 }{ 7 }\]

OpenStudy (anonymous):

so it is basically: (# - mean)^2 + (# - mean)^2 ..... / the total number of #s

OpenStudy (anonymous):

Are you with me @TammisaurusRex ?

OpenStudy (anonymous):

Oh! Okay! Calculating it now (:

OpenStudy (anonymous):

Okaii :)

OpenStudy (anonymous):

It would be 25.14 (rounded), right?

OpenStudy (anonymous):

\[\frac{ 176 }{ 7 } \approx 25.14\]

OpenStudy (anonymous):

So yes, indeed :) Good Job :) So now to find the standard deviation :)

OpenStudy (anonymous):

Oh, I didn't know there was further solving.. my answer options have that double ~ symbol and numbers, I don't think they wanted me to solve it all the way

OpenStudy (anonymous):

Find the Standard Deviation: \[\sigma = \sqrt{variance}\] \[\sigma = \sqrt{\frac{ 176 }{ 7 }}\] \[\sigma = \sqrt{25.14}\] \[\sigma \approx 5.01\]

OpenStudy (anonymous):

\[\approx \] this is just approximately

OpenStudy (anonymous):

Oh! I see (: Thank you!

OpenStudy (anonymous):

yw :) No problem :D